Systems Staff Meeting Notes 3/29/99

Systems Staff Meeting Notes
March 29, 1999

Present: Rich Bennett, Brian Burton, Bill Covey, Debra Harris, Bobbie Meng, Everett Phillips, Suzy Shaw, Ying Tang
Review Minutes of 3/22/99
  1. Ad Hoc #3 - clarification on recycled machines: Systems will need to quick format all machines that are returned us for re-allocation. A student could be trained to do this, but it is necessary due to software license compliance.
Ad Hoc
  1. Security II - review of our departmental lockup checklist - the procedure to be followed by the last person to leave the office for the day. Don't forget the copier & lights.
  2. Server security - NT servers are always to be locked with an administrative password. Do not turn off the servers.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Gateway Orders - it is uncertain in what form the next shipment will be received. We have been in confusing correspondence with Gateway - they had changed the case from tower to desktop on the last order and we've been trying to convert it. Also, they told us that they had "fixed" the ethernet card problem, but we thought they had already canned that request. We are trying to sort it out, but any shipment that comes in the next week or so will probably not have the specifications changed from the last order - that is, they may contain 2 ethernet cards and a desktop case.
Liaisons
  1. Melissa Virus - brought down quite a few sites, today is expected to be the worst. If you receive the message, do not open the attachment. The McAfee site has a patch available. Since it is a macro virus, it will probably get cloned.
  2. Glenda put notes on everyone's monitor this morning in Accounting to warn them of the virus. Such a diligent liaison!
Networking
  1. There was a dial-up outage on Sunday, but was quickly fixed.
  2. CISCO has shipped a replacement box for the Internet router. NERDC is not sure that they want to make the change. They suspect that it is a hardware problem, but want to try one more software change before replacing the box. We are still not on regular configuration - the overall problem has not been resolved yet.
  3. We will be putting out Linux servers soon: 'Squid' will be a web browser used to control web email filtering. If it works, we will also use it for ATC and print queues. We will build them on a Pentium platform. This should provide better local security and backup web proxy access. If it seems to work well, we will put two in each closet, one to be used as backup.
Software Projects Status
  1. The Upload program for Resource Services to batch vendor info into NOTIS is in beta testing with Resource Services.
  2. ATC Standalone - the first version is ready for alpha deployment.
  3. LASSIE - dummy records have been deleted. Waiting for the first order, will input & print out. If all goes well, Maria (from Resource Services) could take care of licensing approvals.
  4. IR database - Ying is finished with the software part and is currently testing it. She will try to get together with Bobbie this week to get her input on the current version.
  5. Mellon - still no word from Harvard.
  6. Order Request Clone (from LASSIE) - no start date yet - need to research and confer with management.
DP Coordinator
  1. Compustat is still hanging.
  2. Next week is Student Employment Recognition Week. Bill has generously offered to recognize our students with the purchase of two movie tickets to AMC theaters for each student.
Calendar
  1. Ashley is out sick today.
  2. Winston will be out March 29 - April 2.
  3. Bobbie out March 31 & April 1.
  4. Everett will be out April 2 and April 5.
  5. Rich will be out April 6-9.
  6. April 5-9 - National Student Employment Week


Last updated February 18, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu